3. Assume that the short-run cost and demand data given in the table below confront a monopolistic competitor selling a given product and engaged in a given amount of product promotion. Compute the marginal cost and marginal revenue of each unit of output and enter these figures in the table.

Output Total cost Marginal cost Quantity demanded
Price Marginal revenue
0 $ 75 0 $180
1 120 45 1 165 $_____
2 135 15 2 150 _____
3 165 30 3 135 _____
4 210 45 4 120 _____
5 270 60 5 105 _____
6 345 75 6 90 _____
7 435 90 7 75 _____
8 540 105 8 60 _____
9 660 120 9 45 _____
10 795 135 10 30 _____

(a) At what output level and at what price will the firm produce in the short run? What will be the total profit?



(b) What will happen to demand, price, and profit in the long run?
1
Expert's answer
2016-04-23T06:47:05-0400
Q TC MC Qd P TR MR
0 $75 - 0 $180 $0 -
1 120 45 1 165 165 $165
2 135 15 2 150 300 135
3 165 30 3 135 405 105
4 210 45 4 120 480 75
5 270 60 5 105 525 45
6 345 75 6 90 540 15
7 435 90 7 75 525 -15
8 540 105 8 60 480 -45
9 660 120 9 45 405 -75
10 795 135 10 30 300 -105

(a) The firm will produce in the short run at such output level, for which MR = MC, so Q = 4 units will be produced (MR = MC between 4 and 5 units produced), and the price will be P = $120. Total profit will be TP = TR - TC = 480 - 210 = $270.
(b) In the long-run new firms will enter the market, because the industry is profitable, so the quantity demanded will increase, the price will decrease, and the firm will earn normal (zero) profit.